OpenStudy (anonymous):

A coin bank contains 250 coins, all dimes and quarters. Together the coins total $39.95. How many quarters are in the coin bank? How many dimes are in the coin bank? I do know i have to use the numbers in an equation, making quarters X and dimes Y. I think I set it up like this... x+y = 39.25 and x+y= 250....i think idk :) Help me please <3

OpenStudy (anonymous):

Second equation is right, but for the first one its cost...a quarter is .25 and dime is .10 .25x+.10y=39.25

OpenStudy (anonymous):

so would i graph, substitute, or do combination???

OpenStudy (anonymous):

Any, but substitution would probably be easiest. You can solve for y in the second equation easy y=250-x and plug that into the other

OpenStudy (anonymous):

the do i plug y into the first equation?

OpenStudy (anonymous):

@CalebBeavers

OpenStudy (anonymous):

Yep so youve got .25x+.10(250-x)=39.25 now can you solve for x?

OpenStudy (anonymous):

yeah thanks :)

OpenStudy (anonymous):

Yeah i just noticed it only wants the number of dimes, so once you have x which is the number of quarters just plug it into x+y=250 to get number of dimes

OpenStudy (anonymous):

okay will do :) thanks again!! lol